LHC 2012 - latest news

Week 14: Machine coordinators :Gianluigi Arduini, Barbara Holzer

Beam Commissioning Planning

 

Date

Summary

Link to report from morning meeting

Monday 9th April
  • 4:50 Ramping down ALICE solenoid and dipole (+compensators)
  • 7:32 beams dumped during squeeze by BLM at TCLA.D6L7 RS 1.3 sec (losses on B2).
  • next: 624 bunches (6 hours stable beams with ALICE off) till Tuesday morning
9:00 meeting
Sunday 8th April
  • 0:48 5:09 Fill 2488 stable beams (264 bunches). Changing luminosity at ATLAS and CMS by separating the beams (1st hour 75%, 2nd hour 100%, 3rd hour 75%, 4th hour 100%). Average intensity at start of ramp: 1.34x1011 ppb and at start of stable beams: 1.31x1011 ppb. Initial maximum luminosity: 9.6x1032cm-2s-1, total delivered luminosity: 10.5 pb-1.
  • 7:17 Fill 2489 stable beams (264 bunches). Changing luminosity at ATLAS and CMS by separating the beams (1st hour 95%, 2nd hour 100%, 3rd hour 95%, 4th hour 100%)
  • 14:54 16:10 Fill 2490 (264 bunches) stable beams. Average intensity 1.5e11 per bunch. Initial luminosity: 11e32cm-2s-1. Losses not higher than 10-15% of dump thresholds all along the cycle (below 10% on RS 1.3s during collisions BP). 5 pb-1delivered luminosity. 
  • 18:20 4:37 Fill 2491 stable beams (624 bunches). Average intensity 1.4e11 ppb. Initial luminosity: 24e32cm-2s-1. 60 pb-1delivered luminosity.
9:00 meeting
Saturday 7th April
  • 02:00 Fill #2482 STABLE BEAMS (1.3x1011 p/b) and 264 bunches with initial luminosity of >0.9x1033 cm-2 s-1.
  • 05:30 End of fill #2482
  • 08:00 Problem with the orbit feedback - investigating
  • 12:13 Fill 2485 (3 bunches test fill) dumped by another OFSU problem during squeeze
  • Fill 2485: Corrected the beam process for going to LHCb tilted crossing without having head-on collisions at LHCb. Tested successfully.
  • Loss maps: B1 H + V, B2 H +  V + longitudinal (-500 Hz)
  • 19:17 20:47 stable beams fill 2486 (84 bunches). Average intensity at start of ramp: 1.35x1011 ppb  and at start of stable beams: 1.33x1011 ppb. Initial luminosity: 2.9x1032cm-2s-1. 1.4 pb-1delivered in 1:36 hour.
  • 22:00 - 23:23 no beam from SPS: trip of main power supply
9:00 meeting
Friday 6th April
  • 00:18 End of fill #2472. 3.5 pb-1 in 4:24 hours (OP dump)
  • 01:00 – 01:50: no beam form LINAC2
  • 01:50-07:30 Damper setting-up for 1.4x1011 p/bunch
  • 07:30-10:30 Asynchronous damp test.
  • FIDEL update. Now chromaticity decay corrected.
  • 12:13 Beam dump due to problem in the storage of the reference for the orbit feedback.
  • Issues with RF cogging
  • 15:00 STABLE BEAM #2479. Initial luminosity 3.5x1032cm-2s-1
  • 17:45 end of fill #2479: 2.9 pb-1 in 3:45 hours
  • 19:00-00:00 Fills (2480 and 2481) with 264 bunches/beam and bunch intensity close to 1.5x1011 p/b. Work on losses in the squeeze and stabilization.
9:00 meeting
Thursday 5th April
  • 00:38 2:04 first stable beams 4 TeV (3 on 3 bunches) (#2469) Initial luminosity 7.8x1030cm-2s-1
  • 02:04 End of fill#2469. 39 nb-1 in 1:50 hours.
  • 5:40 - 9:41 Stable Beams 47 bunches (Atlas low pile-up, Alice satellite-main collisions; 11 indiv bunches 8-9E10 and 3 trains of 12 nominal bunches). Initial luminosity 8.1x1031cm-2s-1. Dumped by CMS BCM.
  • 09:41: End of fill #2470. 1100 nb-1 in 4 hours
  • Problems with the following injection due to to intermittent loss of communication with the FGC. Traced back to an issue with a client of the FGC acquisition (Luminosity scan application).
  • 15:27 STABLE BEAMS #2471 (47 on 47 bunches). Initial luminosity: 8.7x1031cm-2s-1
  • 17:47 End of Fill #2471. 540 nb-1 in ~2:15 hours
  • 19:54 Stable beam #2472. (84 on 84 bunches): Initial luminosity: 2.9x1032cm-2s-1
8:30 meeting
Wednesday 4th April
  • 00:00-04:00 BI setting-up for 50 ns
  • 04:11 Trip of Q4.R8. MP3 piquet called. Pre-cycle.
  • 08:30 Access for QPS for Q4.R8 for reset of QPS card.
  • 10:30-11:40 No beam from the injectors due to unplanned intervention on RAMSES software and SPS extraction septum problem (beam 2)
  • Ramp-Squeeze for chromaticity measurement and verification of the beta*-energy limits. Chromaticity correction during the ramp very good. Still to be improved at flat-top.
  • Energy interlock on the TCTs at IP8 during the squeeze. Corrected.
  • Ramp with new parameters for decayy and snap-back correction (FIDEL) and aperture checks with squeezed optics: problems with the FIDEL parameters, losses at the beginning of the ramp
  • Revert back to previous FIDEL correction version.
  • 14:28 Spurious trigger of the unmaskable BLMs in point 2
  • 16:00 Spurious trigger of CMS BCM ==> Expert contacted.
  • 16:30-21:30: Aperture verification at the end of squeeze and in collision. Aperture sufficient to run with 60 cm optics.
  • 23:00 Injecting.
  • Next: Stable Beams with 3 on 3 bunches
8:30 meeting
Tuesday 3rd April
  • 08:00 Loss maps + asynchronous dump tests at injection
  • Ramp and squeeze with probe beams for chromaticity correction (including setting-up of the b3 decay correction at flat-top). Beam 2 lost in the squeeze from 1 m. Reason: radial modulation and loss of the lock of the phase loop
  • 12:07 Trip of Q4.R8 due to noise on the HTS quench protection system.
  • 15:00 Asynchronous dump tests at injection
  • 15:30-24:00: Setting-up of the 50 ns injection: up to 144 bunches/pulse injected. Some vacuum activity observable at MKI and TDI areas
8:30 meeting

Monday 2nd April

  • 06:00 Dump qualification test at 2 TeV.
  • 08:00 Access to prepare for stable beams (PC, BBQ Pt 4, cryo Pt 2 insulation vacuum, BLM Pt 2+8, experiments)
  • 13:00 machine closed, RQTL10.R3B2 tripped, pre-cycle
  • 14:30 RF blow-up test ramp with new blow-up functions
  • Test of the b3 knob for chromaticity decay correction
  • Loss maps at the end of the squeeze (betatronic and asynch dump test)
  • 20:30 Ramp with old-blow-up. Still some issues with the control from the sequencer
  • Loss maps + asynchronous dump test in collision over night

8:30 meeting